OpenStudy (anonymous):

The length of a rectangle is 2 more than 3 times the width. If the perimeter is 100 meters, what is the width of the rectangle? 13 meters 12 meters 11 meters

OpenStudy (a1234):

2w + (3w + 2)(2) = 100

OpenStudy (anonymous):

i tried that but got a different answer @a1234

OpenStudy (a1234):

Let's see... 2w + (3w + 2)(2) = 100 2w + 6w + 4 = 100 8w = 96 w = 12 So then the length is 38. 76 + 24 = 100 It works :)

OpenStudy (anonymous):

ya but 38 inst a option see that the part that confuses me

OpenStudy (anonymous):

oh 12 oh nvm i feel so stupid im sorry

OpenStudy (a1234):

They're asking you for the width, which is 12 meters.

OpenStudy (a1234):

It's okay.

OpenStudy (anonymous):

thank you

OpenStudy (a1234):

You're welcome.